OpenStudy (mr.potatoman):

What are the coordinates of the centroid of a triangle with vertices D(0, 1) E(2, 6) , and F(7, 2) ?

563blackghost (563blackghost):

http://www.mathopenref.com/coordcentroid.html

563blackghost (563blackghost):

Use the formula provided. \(\huge\bf{O_{x}=\frac{0+2+7}{3}}\) \(\huge\bf{O_{y}=\frac{1+6+2}{3}}\)

563blackghost (563blackghost):

\(\huge\bf{(O_{x},O_{y})=centroid~coordinate }\)

likeabossssssss (likeabossssssss):

@563blackghost nice latex

563blackghost (563blackghost):

Oh thank you :)

likeabossssssss (likeabossssssss):

np:)